OpenStudy (anonymous):

Let f(x)=2x+3 and g(x)-x^2-x find the value of (f/g)(2)

OpenStudy (anonymous):

(f/g)(2)= (2(2)+3)/(2^2-2) =7/2

OpenStudy (anonymous):

ty